OpenStudy (anonymous):

two bodies are thrown vertically upwards with their initial velocity in ratio 2:3 the ratio of maximum height attained will be?

OpenStudy (ajprincess):

say the velocity of the first particle is u1. say the velocity of the second particle is u2. u1/u2=2/3 u1=2u2/3 The velocity at the maximum height will be 0. 0=u1^2-2gh1 u1^2=2gh1 u2^2=2gh2